OpenStudy (anonymous):

Find the acceleration of this particle at all time values where this particle stops

OpenStudy (anonymous):

This should be in physics, but can I have the values please (:

OpenStudy (anonymous):

A particle is moving along the x-axis. At time t, the particle is at x(t)= 2t - 9t^2 +12t +2. Find the acceleration of this particle at all time values where this particle stops.

OpenStudy (anonymous):

I guess you can start by finding the roots of the equation to get you the graph of the function. If the particles stops, there is no acceleration and velocity = 0.

OpenStudy (anonymous):

Hmmmm. We are going over derivatives and differentiation in class, if that makes it make any more sense.

OpenStudy (anonymous):

hmm this must be calculus :P I dont have it till next semester. but ill learn what I can (:

OpenStudy (anonymous):

as of now

OpenStudy (anonymous):

lol, it is indeed calculus, and a pain in my butt! :p

OpenStudy (anonymous):

lool you must be a senior, looking at these online notes make my head spin. It is pretty confusing eh?

OpenStudy (anonymous):

This is what I'm looking at: http://mathworld.wolfram.com/Derivative.html

OpenStudy (anonymous):

that comic strip at the end of the page :P

OpenStudy (anonymous):

hahaha the comic strip was funny and sooooo true! and actually I'm a junior in college hahaha.

OpenStudy (anonymous):

ahah yeah :D and aha oh really? that's neat! Well i'll call for help, see if anyone sees this message, if only i knew this stuff :(

OpenStudy (anonymous):

thanks! I'm meeting a friend now to see if he can help me out! thanks for trying lol.

OpenStudy (anonymous):

lol alrii, it's night time here, so ill just meet my bedd... Night ;P

OpenStudy (anonymous):

@sauravshakya @ujjwal @timo86m @mukushla @TuringTest Not my question, but hai if anyone of you could solve it, that would be great

OpenStudy (anonymous):

find x'(t)

OpenStudy (anonymous):

how would you do that? find roots?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

I think its x(t)= 2t^3 - 9t^2 +12t +2. Now, velocity=x'(t)=6t^2 -18t+12 And accleration=x''(t)=12t Now, when the particle stops , velocity=0 6t^2 -18t+12=0 t^2-3t+2=0 t^2 -2t -t +2=0 t(t-2)-1(t-2)=0 (t-1) (t-2)=0 Thus, t=1 or 2 Now, when t=1 , accleration=12*1=12 when t=2 then , accleration =12*2=24

OpenStudy (anonymous):

does it help?